MICQ(1) MICQ(1) NOME mICQ−unpiccoloclientICQtestuale SINTASSI micq[-h][-c][-i<locale>][-b<basedir>][-f<configfile>][-v[level]][-l<logplace>] DESCRIZIONE mICQe` unclienttestualeperilprotocolloICQchesupportamoltedellecaratteristichefornite dalclientdiMirabilis.mICQpermettedicambiarelapassworddell’utente,crearenuoviaccount, aggiornareleinformazionidell’utente,cercarealtriutenti,inviareSMS,aprireconnessionidirette, inviaremessaggidiconsenso,trasferirefiles,scoprirequaliclientusanoglialtriutenti,inviaremessaggicodificaticonUTF-8,eseguirequalsivogliacomandoperqualsisasievento,ecc. OPZIONI −h|−−help Mostraunbrevetestodiaiutoedesce. −c|−−nocolor Disabilital’utilizzodeicolori. −i|−−i18n<locale> Utilizzalatraduzionelocaleinvecediquelladerivatadallevariabilid’ambiente. −b|−−basedir<basedir> SpecificabasedircomeladirectorydibasedovemICQcerchera` isuoifile,inparticolareil suofilemicqrc. −f|−−config<configfile> Specificaconfigfilecomeilfilediconfigurazionedausare. −v|−−verbose[level] Impostalamodalita` prolissa,oimpostalaprolissitalgradolevel. −l|−−logplacefile|dir Utilizzafilecomelogfileperscrivercileregistrazioni,outilizzadircomedirectoryper crearelogfileperciascunutente. CARATTERISTICHE mICQe` portabilissimoegirasuLinux,DOS,OS/2,Win32,BeOS,AmigaOS,GNUHURDe moltialtrisistemiunixmoderni.E’risaputochelaversione0.4.10dimICQgirisuLinux, AIX,HPUX,Tru64,AmigaOS,parzialmentesuBeOSealtrisistemiunix.Tuttavia,glialtri adattamenti(ports)potrebberofunzionareoppurenoechiunquevogliacollaudarlie` ilbenvenuto. STORIA 16Ago1998 MatthewD.Smithrendepubblicomicqseguitoda12uscitedatatefinoal25Settembre 1998. 05Gen2001 MatthewD.Smithmuoreinunincidenteautomobilistico.Gliutentiproseguonoillavoro diMatthewdadoveeglilolascio` ,formalizzanomICQsottolalicenzaBSDeriniziarono losviluppodimICQ. 10Giu2002 Laversione0.4.9vienerilasciatasottolalicenzaGPL,implementandoinfinelanuovaversione8delprotocolloICQ,implementazionefattadaRu¨ digerKuhlmann. MAILING-LIST LamailinglistdimICQe` micq-list@micq.org.Leinformazionicircalasottoscrizioneel’archivio sonodisponibiliahttp://www.micq.org/pipermail/micq-list/. DOWNLOAD ViaHTTPdahttp://www.micq.org/download. Perottenerelaversionepiu` nuovaconilsistemaCVS,digitare: IT mICQ 1 MICQ(1) MICQ(1) cvs−d:pserver:anonymous@micq.org:/var/lib/cvscomicq VEDIANCHE micqrc(5),micq(7) FILE ˜ /.micqrc VecchiofilediconfigurazionepermICQ−obsoleto. ˜ /.micq/ LadirectorydibasepredefinitadovemICQricercaisuoifile. BASE/micqrc FilediconfigurazionepermICQ. BASE/i18n/*.i18n FileditraduzionedimICQperignorareleimpostazionilocali. BASE/contacts/ Directorydoveimetadatipericontattisonoconservati. BASE/history/ Directorypredefinitaperifilediregistro(logfile). BASE/files/ Directorydoveifileinrecezionevengonosalvati. BASE/remote-control LocazionepredefinitadelFIFOusatoperilcontrolloremoto. PREFISSO/share/micq/*.i18n FilediinternazionalizzazionecontenentelatraduzionedituttelestringhechemICQ mostra. DIFETTI(BUGS) Inviairesocontideidifetti(bugreports)ocommentiaRu¨ digerKuhlmann<micq@ruediger-kuhlmann.de>(inTedescooinInglese). AUTORI L’autorediquestapaginadimanualee` JamesMorrison<ja2morrison@student.math.uwaterloo.ca> ede` derivatadallapaginadimanualeminimalescrittadaJordiMallach<jordi@sindominio.net> perilsistemaDebianGNU/Linux.Questapaginadimanualee` statatradottainitalianoda FabioVenturi<fabio.venturi@studenti.unipr.it>. IT mICQ 2